### v1

@@ 1 Introduction
A variety of methods have been employed to study GPP. In some cases, it is possible to approximate particle production using various analytical methods, but for a full analysis, numerical methods must be employed. In order to understand the spectra of produced particles, it is thus important to fully understand where each analytic method is valid and how these approximations compare to the full numerical calculation. The relevant analytic methods are as follows: first, one can trace out the evolution of each mode as it transitions from the inflationary quasi-de Sitter regime into matter and/or radiation domination by solving the mode equation in each region and matching at the boundaries. This method we refer to as ‘boundary matching.’ One can also employ the Stokes phenomenon Hashiba:2021npn and steepest descent PhysRevD.67.083514 methods to approximate GPP. Finally, GPP can be analytically approximated by considering inflaton scattering processes, as discussed in e.g. Tang_2017 ; Basso_2022 . In this work, we perform a detailed analysis of the GPP for scalars, considering both analytic and numerical methods for the GPP of both conformally and minimally-coupled scalars. For the background inflationary dynamics we consider quadratic inflation and both early and late reheating scenarios. We analytically approximate the GPP in each of these cases and outline the regime of validity of each analytic method. We additionally numerically calculate the GPP and compare with the analytic results, finding good agreement between the two.
@@ 2 Inflation and Gravitational Particle Production
This evolution is conveniently engineered by assuming that there is a single scalar field - the inflaton field $\varphi$, whose potential energy drives an inflationary phase with $\ddot{a}>0$ (dot is derivative with respect to coordinate time). Inflation ends when the kinetic energy of the inflaton field dominates the potential energy and $\ddot{a}$ first becomes negative. The inflaton field then undergoes coherent oscillations about the minimum of its potential, and its energy density redshifts as matter ($\rho_{\varphi}\propto a^{-2/3}$) when averaged over several oscillations. During this MD era, the energy density evolves as $\exp{(-\Gamma_{\varphi}t)}$, where the decay width, $\Gamma_{\varphi}$, is the rate at which the inflaton decays into relativistic standard model particles. Eventually, the radiation dominates the inflaton energy density, at which point the Universe “reheats” and enters the RD epoch.
We consider two reheating scenarios: early reheating and late reheating. Late reheating occurs when $(a_{\mbox{\tiny RH}}/a_{e})>(m/H_{e})^{-2/3}$, i.e., when $H_{\mbox{\tiny RH}}<m$. In numerical calculations, late reheating is implemented by setting $\Gamma_{\varphi}=0$. For early reheating, there exists a limiting case of “immediate reheating,” where reheating occurs instantaneously at the end of inflation.
@@
The production of neutrinos during nucleosynthesis places a lower bound of $T_{{\mbox{\tiny RH}}}^{\rm min}\approx 5$ de_Salas_2015 . The maximum $T_{{\mbox{\tiny RH}}}$ depends the energy scale of inflation, as
@@ 3.2 Minimal coupling
We now again find the comoving number density $\tilde{a}^{3}\tilde{n}$ from $\tilde{a}^{3}\tilde{n}_{k}$. Unlike for conformally-coupled scalars where the spectra are blue-tilted before reaching ${\tilde{k}}_{\mathrm{peak}}$, the spectra of minimally-coupled scalars are red-tilted in the low-${\tilde{k}}$ region, and continue to decrease as ${\tilde{k}}$ increases. This means that the $\tilde{k}$ range we have calculated numerically only captures the “tail end” of the area under the curve, and modes with smaller ${\tilde{k}}$ carry more of the number density. Therefore, before finding $\tilde{a}^{3}\tilde{n}$ by integrating the ${\tilde{a}^{3}\tilde{n}_{k}}$ spectrum, we first extrapolate the spectrum to ${\tilde{k}}=10^{-20}$ by fitting a line to the low-${\tilde{k}}$ region. This value of minimum ${\tilde{k}}$ is chosen since CMB observations probe scales of approximately $10^{4}{\tilde{k}}_{0}$, where ${\tilde{k}}_{0}=a_{0}H_{0}/a_{e}H_{e}$ are the modes that are of a scale equal to the Hubble scale today. Using $a_{0}/a_{\mbox{\tiny RH}}=T_{\mbox{\tiny RH}}/T_{0}$ and $a_{\mbox{\tiny RH}}/a_{e}=(H_{e}/H_{\mbox{\tiny RH}})^{3/2}$, we find ${\tilde{k}}_{0}\sim 10^{-24}(10^{9}{\mathrm{GeV}}/T_{\mbox{\tiny RH}})^{1/3}(10^{12}{\mathrm{GeV}}/H_{e})^{1/3}$. Therefore, the spectrum is extended to ${\tilde{k}}=10^{4}{\tilde{k}}_{0}=10^{-20}$.

### v2

@@ 1 Introduction
A variety of methods have been employed to study GPP. In some cases, it is possible to approximate particle production using various analytical methods, but for a full analysis, numerical methods must be employed. In order to understand the spectra of produced particles, it is thus important to fully understand where each analytic method is valid and how these approximations compare to the full numerical calculation. Previous work in the literature have discussed the aspects of the various analytic methods individually, but without detailed discussion of when they are valid. The relevant analytic methods are as follows: first, one can trace out the evolution of each mode as it transitions from the inflationary quasi-de Sitter regime into matter and/or radiation domination by solving the mode equation in each region and matching at the boundaries. This method we refer to as ‘boundary matching.’ One can also employ the Stokes phenomenon Hashiba:2021npn and steepest descent PhysRevD.67.083514 methods to approximate GPP. Finally, GPP can be analytically approximated by considering inflaton scattering processes, as discussed in e.g. Tang_2017 ; Basso_2022 . In this work, we perform a detailed analysis of the GPP for scalars, considering both analytic and numerical methods for the GPP of both conformally and minimally-coupled scalars. For the background inflationary dynamics we consider quadratic inflation and both early and late reheating scenarios, the former of which do not yet appear in the literature. We analytically approximate the GPP in each of these cases, and outline the regime of validity of each analytic method. We additionally numerically calculate the GPP and compare with the analytic results, finding good agreement between the two. We aim to present a comprehensive picture of the analytic and numerical methods for calculating GPP, with a particular focus on the full spectrum of analytic approximations.
@@ 2 Inflation and Gravitational Particle Production
This evolution is conveniently engineered by assuming that there is a single scalar field - the inflaton field $\varphi$, whose potential energy drives an inflationary phase with $\ddot{a}>0$ (dot is derivative with respect to coordinate time). Inflation ends when the kinetic energy of the inflaton field dominates the potential energy and $\ddot{a}$ first becomes negative. The inflaton field then undergoes coherent oscillations about the minimum of its potential, and its energy density redshifts as matter ($\rho_{\varphi}\propto a^{-3}$) when averaged over several oscillations. During this MD era, the energy density evolves as $\exp{(-\Gamma_{\varphi}t)}$, where the decay width, $\Gamma_{\varphi}$, is the rate at which the inflaton decays into relativistic standard model particles. Eventually, the radiation dominates the inflaton energy density, at which point the Universe “reheats” and enters the RD epoch.
We consider two reheating scenarios: early reheating and late reheating. Late reheating occurs when $(a_{\mbox{\tiny RH}}/a_{e})>(m/H_{e})^{-2/3}$, i.e., when $H_{\mbox{\tiny RH}}<m$. In numerical calculations, late reheating is implemented by setting $\Gamma_{\varphi}=0$. This corresponds to a reheat time of $a_{{\mbox{\tiny RH}}}=\infty$; we assume that all of the particle production occurs sufficiently before the reheating epoch such that the particle production arises from the quasi-de Sitter and matter-dominated eras and any details of reheating are irrelevant. For early reheating, there exists a limiting case of “immediate reheating,” where reheating occurs instantaneously at the end of inflation.
@@
The production of neutrinos during nucleosynthesis places a lower bound of $T_{{\mbox{\tiny RH}}}^{\rm min}\approx 5$ MeV de_Salas_2015 . The maximum $T_{{\mbox{\tiny RH}}}$ depends the energy scale of inflation, as
@@ 3.2 Minimal coupling
We now again find the comoving number density $\tilde{a}^{3}\tilde{n}$ from $\tilde{a}^{3}\tilde{n}_{k}$. Unlike for conformally-coupled scalars where the spectra are blue-tilted before reaching ${\tilde{k}}_{\mathrm{peak}}$, the spectra of minimally-coupled scalars are red-tilted in the low-${\tilde{k}}$ region, and continue to decrease as ${\tilde{k}}$ increases. This means that the $\tilde{k}$ range we have calculated numerically only captures the “tail end” of the area under the curve, and modes with smaller ${\tilde{k}}$ carry more of the number density. Therefore, before finding $\tilde{a}^{3}\tilde{n}$ by integrating the ${\tilde{a}^{3}\tilde{n}_{k}}$ spectrum, we first extrapolate the spectrum to ${\tilde{k}}=10^{-20}$ by fitting a line to the low-${\tilde{k}}$ region. This value of minimum ${\tilde{k}}$ is chosen since CMB observations probe scales of approximately $10^{4}{\tilde{k}}_{0}$, where ${\tilde{k}}_{0}=a_{0}H_{0}/a_{e}H_{e}$ are the modes that are of a scale equal to the Hubble scale today. Using $a_{0}/a_{\mbox{\tiny RH}}=T_{\mbox{\tiny RH}}/T_{0}$ and $a_{\mbox{\tiny RH}}/a_{e}=(H_{e}/H_{\mbox{\tiny RH}})^{3/2}$, we find ${\tilde{k}}_{0}\sim 10^{-24}(10^{9}{\mathrm{GeV}}/T_{\mbox{\tiny RH}})^{1/3}(10^{12}{\mathrm{GeV}}/H_{e})^{1/3}$. Therefore, the spectrum is extended to ${\tilde{k}}=10^{4}{\tilde{k}}_{0}=10^{-20}$. We also note that as a dark matter candidate, light, minimally coupled scalars face cosmological challenges, as they induce large isocurvature fluctuations which are stringently constrained by the CMB PhysRevD.64.023508 . Beyond dark matter, if the produced particles are unstable, they may decay and be unconstrained by CMB isocurvature. Therefore, we include analysis of such models here for completeness.